How do i solve this problem using the thermodynamic structure?
- what will be the temperature change when 75 cal of heat are added to 75 g of water?
1
Expert's answer
2011-01-10T17:46:49-0500
It can be found using the formula:

Δ T = Q /(m*Cm)
Q = 75 cal= 313.8 J
m = 0.075 kg
Cm = 4.18 kJ/kg*grad

Δ T = Q /(m*Cm) = 313.8 / (0.075*4.18x103) = 1.

The temperature will increase by one Kelvin.